# Minerva: A File-Based Ransomware Detector

### v2

@@ 6.7.2. Sample prediction time
Minerva relies on fully-connected DNN classifiers to determine whether a file activity is benign or malicious. To provide an overview of the time overhead associated with Minerva’s prediction process, we invoked the predict functionality of Minerva’s detector on $1,000$ individual samples and recorded the timing. On average, the Minerva’s classifiers take $0.0138sec$ to predict a single sample ($0.0016sec$ standard deviation), introducing negligible impact on the system and Minerva’s decision-making pipeline.
@@ 7. Discussion and Limitations
As outlined in Section 6, Minerva shows significant effectiveness in promptly identifying ransomware activity within the system. Depending on the encryption method adopted by the ransomware, however, there exists a potential for data loss. Our evaluation indicates that, for $K=2$, Minerva achieves an average detection time of $0.52sec$. This corresponds to approximately 19 files encrypted before detection, on average (Splunk, 2022). Nonetheless, certain ransomware families such as LockBit and Babuk demonstrate much higher speeds, capable of encrypting up to ${\sim}130$ files within the same time frame (Splunk, 2022). Furthermore, in theory, ransomware could generate a separate process for each file in the system to execute the encryption process, which would result in complete data loss. While such an approach would be impractical and quickly detectable through alternative means, sophisticated ransomware could employ a simplified version of this encryption method to maximize data loss. We emphasize that this concern is not exclusive to Minerva, but is a shared consideration for all ransomware detection systems.
Owing to its rapid detection rate, Minerva enables the utilization of near-zero overhead data recovery methods to effectively eliminate the risk of data loss. Modern operating systems employ a technique known as I/O buffering, which utilizes a page cache to temporarily store data in main memory before writing it to an I/O device. Operating systems periodically flush portions of the page cache: Windows flushes $1/8th$ of the cache every second (Russinovich et al., 2012), while Linux, by default, holds pages for a maximum of 30 seconds. Given that Minerva’s detection rate largely falls below the cache flushing rate of both operating systems, it becomes feasible to seamlessly adopt cache-based recovery techniques such as (Elkhail et al., 2023). These techniques modify cache flushing policies to prevent content flagged as malicious in the cache pages from being written to disk, providing a near-zero overhead data recovery process.
@@
Lastly, another limitation of Minerva lies in its coarse-grained detection. Minerva detects ransomware activity on a per-file, per-window basis. While this allows Minerva to detect complex attacks such as evasive multiprocess ransomware, it also limits the detection granularity at the file, rather than at the process level. Thus, when Minerva detects ransomware activity on a file during a given time window, all processes that were acting on the file during the window are classified as ransomware and suspended. This group may include benign processes, that wrote data on the file within the same window as the ransomware. We argue, however, that this is a minor limitation. First, operating systems prevent concurrent writes on the same file. Therefore, while the ransomware is encrypting a file, no benign write activity can happen. Second, since Minerva detection time is very low, even if benign activity happened in the same window as ransomware, on average, only $0.5sec$ of benign activity would be lost. Lastly, we argue that the benefits of promptly stopping ransomware activity, outweigh the disadvantage of dropping a very short amount of benign user activity.
@@ 8. Conclusions
9. Acknowledgments
This work was supported by the PON program of the Italian MUR under the project “Application of Machine Learning to improve olive yield and reduce climate change impact”, project SERICS (PE00000014) under the NRRP MUR program funded by the EU-NextGenerationEU, and project AutoAD: Using Active Defense to Defeat Cyber Adversaries funded by Sapienza University of Rome.

### v3

@@ 6.6.2. Sample prediction time
Minerva’s detectors leverage fully-connected DNN classifiers to identify ransomware activity. We evaluate the overhead of Minerva’s Behavioral Detector module in predicting file activity by calling the predict function on $1,000$ individual samples sequentially. The prediction function invokes the individual prediction functions of each tier’s classifier, characterizing the overall time required to make a prediction across the full multi-tier architecture of Minerva. Our tests indicate that Minerva’s classification requires on average $0.0138(\pm 0.0016)sec$ to predict a sample, introducing negligible impact on the system and Minerva’s decision-making pipeline. Furthermore, while we tested sequential classification to evaluate the worst-case scenario, we stress that, in practice, classification operations can be parallelized in large batches thanks to the small size of Minerva’s feature vectors (seven float32 features, or 224 bits each).
@@ 7. Discussion and Limitations
We identify three main limitations of Minerva: (1) potential data loss before detection, (2) coarse-grained detection granularity, and (3) potential susceptibility to advanced evasive ransomware.
Data Loss. While Minerva demonstrates significant effectiveness in promptly identifying ransomware activity, there remains a potential risk of data loss before detection occurs. A detection time of $0.52(\pm 0.25)$ seconds, as achieved by Minerva, translates to an average of 19 files being encrypted before detection by contemporary high-speed ransomware (Splunk, 2022). Advanced ransomware families such as LockBit and Babuk are capable of encrypting up to ${\sim}130$ files within the same time frame (Splunk, 2022), and future ransomware families will certainly keep improving. While currently Minerva does not implement file recovery techniques, these can be implemented in the future to prevent any data loss. Given Minerva’s rapid detection rate, transparent, near-zero overhead data recovery methods such as cache-based backups can be utilized (Elkhail et al., 2023). Modern operating systems heavily rely on I/O buffering, which utilizes a page cache to temporarily store data in the main memory before transferring it to I/O devices. Portions of the page cache are periodically flushed at different intervals depending on the operating system: $1/8th$ of the cache every second for Windows (Russinovich et al., 2012), while Linux retains pages for up to 30 seconds by default. Since Minerva’s detection rate is significantly faster than the cache flushing rate, it can be integrated with techniques that adjust cache flushing policies to prevent flagged malicious content from being written to disk. This approach would enable a near-zero overhead data recovery process (Elkhail et al., 2023).
@@
Detection Granularity. Minerva detects ransomware activity on a per-file, per-window basis. This behavior inherently limits the detection granularity to the file level rather than the process level. When Minerva detects ransomware activity on a file in a time window, all processes that interacted with the file during that window are classified as ransomware and suspended. This approach may include benign processes that wrote data to the file within the same window as the ransomware. We argue, however, that this is a minor limitation. First, operating systems prevent concurrent writes on the same file. Therefore, while the ransomware is encrypting a file, no benign write activity can happen. Second, since Minerva detection time is very low, even if benign activity happened in the same window as ransomware, on average, only $0.5sec$ of benign activity would be lost. While this might be acceptable for many types of usage, it might not be suitable in high-frequency I/O environments. A partial solution to this problem is to integrate process control and process analysis tools into Minerva. When ransomware activity is detected on a file, Minerva would freeze all processes interacting with it. Simultaneously, the user would receive a notification about the malicious activity and be prompted to confirm whether each identified process is legitimately accessing the file. Unverified processes would be terminated, and their file operations could be rolled back using the previously discussed cache-based recovery method. Such an approach would require the implementation of tools to help users identify processes and cannot be fully automated reliably. However, we argue that the benefits of promptly stopping ransomware activity outweigh the potential disadvantages.
Advanced Evasive Ransomware. As ransomware evolves, future variants may adopt increasingly sophisticated evasion techniques. While Minerva is robust to several types of evasion attacks, a dedicated adversary may be able to still evade detection. For instance, ransomware could be designed to generate separate encryption processes for each file in the system, which would lead to high data loss. While creating so many new processes would be quickly detectable, sophisticated ransomware could employ a simplified version of this encryption method to maximize data loss. Advanced ransomware could coordinate its file activity with that of benign processes, reducing the effectiveness of some of Minerva’s features such as read/write entropy, and potentially leading to evasion. Finally, given the reliance of Minerva on DNNs, future ransomware could utilize adversarial attacks leveraging carefully crafted noise to achieve evasion, such as adversarial examples (Grosse et al., 2017). While such attacks are complex and require advanced knowledge, they are not impossible and could lead to failed detection. Extending Minerva to consider further complex, coordinated adversarial attacks remains an open research direction.
@@ 8. Conclusions
Appendix A Additional Setup Details
A.1. Dataset Analysis
Data presented in Figure 2 is computed over benign, traditional ransomware, and evasive ransomware process operations performed over a one-second time window. To improve the presentation, we filtered the dataset by removing null read and write operations (which are also ignored by Minerva).
A.2. Dataset description
The dataset used in our experiments comprises data from multiple sources (Continella et al., 2016; Mehnaz et al., 2018; De Gaspari et al., 2022a; Zhou et al., 2023) and additional data gathered as part of this study. Table 4 lists the types of ransomware, ransomware families, and the number of samples for each family. The dataset includes five Traditional Classic ransomware families for a total of 383 samples, which are gathered from previous works (Continella et al., 2016). We additionally gathered 43 samples from five different ransomware families released between 2023 and 2025, referred to as Traditional Current in the table. We use 30 evasive multiprocess ransomware samples (De Gaspari et al., 2022a), and three adaptive ransomware families designed specifically to evade Minerva totaling 20 samples (see Section 5.2 for details on adaptive families). Traditional Classic ransomware samples from (Continella et al., 2016) were gathered in 2016, however, no details are available on when each sample was first seen in the wild. Evasive multiprocess ransomware are research prototypes and have no ransomware families. We use data from all proposed evasive configurations: process splitting, functional splitting, and mimicry (De Gaspari et al., 2020, 2022a). Adaptive multiprocess ransomware are derived from the mimicry evasive multiprocess ransomware (De Gaspari et al., 2022a) and are designed specifically to evade Minerva. We divide them into families based on the group of Minerva features that are manipulated to attempt evasion, as presented in Section 6.4.
Table 5 provides information on the benign dataset. The dataset of benign operations used in this study was provided by (Continella et al., 2016). The dataset was collected from 11 desktop computers over multiple weeks. It comprises computers dedicated to three different types of use: development (dev), office, and home. The dataset includes a wide variety of applications, including office suite applications (.doc, .ppt, .xls), development tools such as Visual Studio, and compression utilities such as WinRar (Continella et al., 2016), providing a varied and challenging setting for behavioral modeling.
Ransomware Year No. Samples
Traditional Classic 383
Crypto Wall 2016 157
Crowti 2016 125
Crypto Defense 2016 77
CTB Locker 2016 14
TeslaCrypt 2016 10
Traditional Current 43
Akira 2023-25 10
LockBit 2023-25 10
Conti 2023-25 10
AvosLocker 2023-25 10
WannaCry 2023-25 3
Evasive Multiprocess / 30
Adaptive 20
Adaptive Entropy / 5
Adaptive Op. Num. / 9
Adaptive R/W Ratio / 6
Total 476
Table 4. Ransomware families used in the evaluation.
Machine Win. Ver. Type Data (GB) IRPs (Mln) Proc. (Mln) Apps Duration (h)
1 10 dev 3.4 230.8 16.6 317 34
2 8.1 home 2.4 132.1 9.67 132 87
3 10 office 0.9 54.2 5.56 225 17
4 7 home 4.7 279.9 18.70 255 122
5 7 home 2.2 138.1 5.04 141 47
6 10 dev 1.8 100.4 10.30 225 35
7 8.1 dev 0.8 49.0 3.28 166 8
8 8.1 home 0.8 43.9 6.33 148 32
9 8.1 home 7.7 501.8 24.20 314 215
10 7 home 0.9 57.6 2.63 151 18
11 7 office 2.6 175.2 4.69 171 28
Total 28.2 1,763.0 107.00 2245 643
Table 5. Benign dataset details.
A.3. Ransomware Data Collection
We collected ransomware data on a Windows 11 virtual machine populated with $\sim 33,000$ files distributed over various subdirectories within the User Windows directory. We included over 80 different types of files, including different image formats (.jpg, .jpeg, .png, .gif), audio/video (.mp3, .mp4, .mov), compressed archives (.zip, .gz), and office files (.docx, .xlsx, .pptx). After each ransomware execution, we verified that the user files were successfully encrypted and discarded samples that did not correctly execute. The VM was reverted and reset after each execution. In total, we gathered 43 working samples from 5 different families, as illustrated in Table 4 (Traditional Current).
Approach Traditional $TPR$ Multiprocess $TPR$
ShieldFS (Continella et al., 2016) 100.00 0.00
RWGuard (Mehnaz et al., 2018) 98.45 0.00
Minerva (ours) 99.25 99.98
A.4. User Study Details
We recruited five volunteer bachelor and master computer science students to carry out our user study. Each student was allowed to freely interact with two Windows systems deployed in a virtual machine: (1) standard system and (2) Minerva system. To ensure that key actions were evaluated, the users were asked to include in their workflow activities typically carried out by different types of users. Participants were asked to include the use of a web browser (standard user activity), edit documents on the local machine (office activity), perform software installation and removal operations (administration activity), and carry out file transfer operations. After interaction with each system, they were asked to rate the system usability on a scale from 1-5, considering in particular whether the system was responsive to interaction and whether activities were unusually slow.
Appendix B Comparison with SOTA
We compare our approach against two state-of-the-art ransomware detection approaches: ShieldFS (Continella et al., 2016) and RWGuard (Mehnaz et al., 2018). As highlighted in recent publications (De Gaspari et al., 2022a; Zhou et al., 2023), ShieldFS and RWGuard still represent reference benchmarks for behavioral-based ransomware detection. Table 6 presents our comparison. As illustrated in the table, Minerva achieves performance comparable to state-of-the-art approaches against traditional ransomware. Furthermore, Minerva successfully detects evasive multiprocess ransomware, a feat that existing approaches fail to accomplish.